Find an equation of the line containing the given pair of points. Use function notation to write the equation.

(2/3 , 4/6) and (-1/3 , 5/6)

Answer by Alan3354(69443)

Find an equation of the line passing through the given points.
============
Same problem, different numbers:
----------
(3,1) and (-2,8)
---
Slope m = diffy/diffx = (8-1)/(-2-3) = -7/5
----
Use y-y1 = m*(x-x1) where (x1,y1) is either point.
y-1 = (-7/5)*(x-3)
5y-5 = -7x + 21
7x + 5y = 26